# Model-guided Fuzzing of Distributed Systems

### v2

@@ Abstract.
We present a coverage-guided testing algorithm for distributed systems implementations. Our main innovation is the use of an abstract formal model of the system that is used to define coverage. Such abstract models are frequently developed in early phases of protocol design and verification but are infrequently used at testing time. We show that guiding random test generation using model coverage can be effective in covering interesting points in the implementation state space. We have implemented a fuzzer for distributed system implementations and abstract models written in TLA+. Our algorithm shows better coverage over purely random exploration as well as random exploration guided by different notions of scheduler coverage and mutation. In particular, we show consistently higher coverage and detect bugs faster on implementations of distributed consensus protocols such as those in Etcd-raft and RedisRaft. Moreover, we discovered 13 previously unknown bugs in their implementations, four of which could only be detected by model-guided fuzzing.
@@
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXccs: Software and its engineering Software testing and debuggingccs: Theory of computation Distributed computing models
@@ 1. Introduction
Coverage-guided fuzzing, which guides test generation toward more coverage, has been effective in exploring diverse executions, mainly in the sequential setting, using structural coverage criteria as a feedback mechanism (Heuse_AFL_2022, ; zeller2019fuzzing, ). However, adopting coverage-guided fuzzing for testing distributed system implementations is nontrivial since there is no common notion of coverage for distributed system executions. Unfortunately, structural code coverage criteria such as line coverage can ignore the orderings of message interactions in a system, thus missing interesting schedules. On the other hand, more detailed criteria, such as traces of messages, may provide too many coverage goals and thus consider each random trace a new behavior, giving up the advantages of coverage-guided exploration.
@@
In this paper, we propose to use state coverage in an abstract formal model of the system as a coverage criterion and present model-guided fuzzing of distributed systems. Abstract formal models are often developed in the design phase of distributed systems to model and formally analyze the underlying protocols (DBLP:books/aw/Lamport2002, ; DBLP:conf/pldi/DesaiGJQRZ13, ; DBLP:conf/pldi/DeligiannisDKLT15, ; DBLP:conf/cloud/DeligiannisGLQ21, ; DBLP:conf/sosp/BornholtJACKMSS21, ; DBLP:journals/cacm/NewcombeRZMBD15, ). We show that these artifacts are also beneficial in the continuous testing infrastructure of the implementations themselves. Our experiments show that a formal model can serve as a good “guide” for a random testing engine—this is because the formal model often captures the important scenarios of the protocol, and coverage of states in the model correlates well with coverage of interesting behaviors in the implementation.
@@
Our motivation for using abstract model guidance to generate test executions shares common insights with semantic fuzzing (DBLP:conf/issta/PadhyeLSPT19, ) and grammar-based fuzzing approaches (DBLP:journals/sigsoft/LePPLVS19, ; DBLP:conf/sigsoft/GopinathMZ20, ) used for testing sequential programs. Semantic fuzzing aims to cover interesting program executions processing program inputs rather than spending exploration budget for exercising uninteresting, syntactic input parsing logic. While a naive fuzzer is likely to generate inputs that cannot pass the input validation and parsing stage, semantic fuzzing generates test inputs that can go deep into the execution. Similarly, a naive event scheduler for distributed systems is likely to produce tests that spend execution budget in exercising uninteresting, network setup stages. For example, it can explore many different orderings of vote messages during the cluster’s leader election phase, barely electing a leader after a prolonged execution. Our approach aims to direct testing toward interesting system behaviors, e.g., processing of user requests once a leader is elected. Similar to grammar-based fuzzing that uses formal specification of the test input to guide test generation, we use an abstract formal model of distributed systems to guide the generation of semantically interesting temporal event schedules.
@@
We have implemented our algorithm for testing distributed systems implementations using TLA+ models (DBLP:books/aw/Lamport2002, ) of protocols. In a nutshell, our testing algorithm proceeds as follows. We start by exploring random schedules of messages, but feed the same sequence of messages to the TLA+ model. We use the TLC model checker (DBLP:conf/charme/YuML99, ) to generate the set of reachable model states corresponding to the explored schedule. We mark a schedule as “interesting” if it covers a new state of the abstract model. We perform, as in coverage-guided fuzzing, a mutation of an interesting schedule by swapping the receipt order of two randomly chosen messages or changing the processes to crash. Applying a mutation to an event schedule gives a new schedule to explore that is similar to the original schedule but likely to exercise new system behavior.
We applied our algorithm to test the implementations Etcd-raft and RedisRaft. Our evaluation shows that model-guided fuzzing leads to higher coverage and can detect bugs faster than pure (unguided) random testing, structural code-coverage guided fuzzing, and trace-based coverage-guided fuzzing. Besides reproducing known bugs, we discovered 13 previously unknown bugs in the implementations of Etcd-raft and RedisRaft. Moreover, four of the new bugs could only be detected by model-guided fuzzing.
@@
    We implement ModelFuzz for testing the two production implementations of the Raft protocol and evaluate it compared to the existing approaches.
@@
    We discovered 13 unknown bugs in total in the implementations of Etcd-raft and RedisRaft.
@@ 2. Overview
Figure 1 presents an example system implemented in a Coyote-like actor programming framework (DBLP:phd/us/Agha85, ; DBLP:conf/cloud/DeligiannisGLQ21, ). The system consists of three processes AppMaster, Worker, and Terminator. Each runs in a separate process and communicates with each other by exchanging asynchronous messages. AppMaster receives client requests, coordinates their execution by a Worker process, and manages the termination of the worker using a Terminator process. It accepts the registration messages from Worker and Terminator and registers them (lines 5-11). Upon receipt of a client request Request(r), it checks whether the cluster is ready by checking the registrations of the worker and terminator. If it is ready, it sends Execute(r) to Worker to handle the request and sends Terminate(w) to Terminator (lines 13-17). The Worker and Terminator register to the AppMaster upon their initialization (lines 26, 39). The Worker handles Execute(r) by processing it (line 30) and Flush by cleaning up its buffers (line 34). The Terminator handles Terminate by sending Flush to the worker (line 42).
@@ 3 StateMachine terminator;
16 terminator.send(worker);
@@ 24 Buffer buffer;
The above implementation has a message race bug (DBLP:conf/asplos/Leesatapornwongsa16, ), which occurs in a particular delivery ordering of the asynchronous messages. The worker code mistakenly omits to check if the buffer is null before processing a task (line 31). This causes the worker to access a null pointer while processing the client request if Flush is delivered to it before Execute(r). Figures 1 and 1 illustrate a correct and an incorrect execution of the implementation.
@@
Coverage-guided testing guides the generation of test executions toward unexplored system behaviors to increase test coverage and search for bugs more efficiently. These methods track the coverage of each test execution and use this information while generating new test cases. However, structural test coverage metrics, such as line or branch coverage, are ineffective for assessing the coverage of distributed system executions since they do not capture the concurrency behavior of the distributed systems. For example, both correct and buggy executions in Figures 1 and 1 hit exactly the same lines and branches of the program in Figure 1 but with different processing orders of the messages.
@@
Instead, we propose to use an abstract, formal, model of the system to assess test coverage and guide the test generation.
@@
[…]
We could have alternatively considered trace-based coverage based on the definition of Mazurkiewicz traces arising out of an execution. Our state-based notion is coarser than traces while keeping the essential information about the system’s behavior. For example, although the messages Register(w) and Register(t) are dependent, their relative ordering does not affect the system state. On the other hand, the ordering of Request(r) affects the reached system state; the system handles the request only if it is delivered after the two registration messages. The given system has 10 possible message orderings with 8 Mazurkiewicz traces (capturing the commutativity of the Execute and Terminate messages). However, the set of all possible system states can be covered by running fewer executions, e.g., only 2 executions for this example.
@@
Suppose the following set of executions has been explored:
@@
Coverage-guided testing selects the executions that hit previously unseen coverage classes and generates new test executions by mutating them. Trace-based coverage would label all these executions as interesting since each belongs to a different coverage class (i.e., they deliver Request(r), Register(w), and Register(t) to the same process in a different order). Generating new tests around all these executions leads to a high number of redundant executions since many of them already produce the same system behavior.
@@ 3. Coverage-Guided Fuzzing with Abstract Models
We use coverage-guided fuzzing for exploring distributed system executions and guiding the test generation using abstract model coverage. Our method is complementary to the traditional fuzzing methods (e.g., AFL (url/afl, )), which explore the space of program inputs, as ModelFuzz explores the space of event schedules in distributed systems. In this section, we define executions and explain how we adopt the coverage-guided fuzzing approach for the model-guided exploration of distributed system executions.
@@ 3.1. Executions of Distributed Systems
A distributed system ${S}$ consists of a set of processes that concurrently operate on their local states and communicate by exchanging asynchronous messages. The processes are equipped with message buffers that keep the messages sent to that process, and they process the messages in their buffers serially. Upon handling a message, a process can update its local state and/or send new messages to the processes.  [^2] We consider FIFO message queues that preserve the order of messages between the same sender-receiver pairs, as in P (DBLP:conf/pldi/DesaiGJQRZ13, ), Coyote (DBLP:conf/cloud/DeligiannisGLQ21, ), or Akka (url/Akka, ).
[…]
@@ 3.2. Coverage-guided Fuzzing of Distributed Systems
Output: The set of covered abstract states $totalCoverage$
@@ 10                         $p\leftarrow\mathtt{assignEnergy}(coverage)$
12                               $T^{\prime}=\mathtt{mutate}(t)$ ; $T\leftarrow T\cup T^{\prime}$
@@ 21       $states\leftarrow\mathtt{executeAbstractModel}(M,actions)$
Algorithm 1 Coverage-guided fuzzing. The statements that differ from traditional fuzzers are highlighted in blue.
Algorithm 1 shows coverage-guided fuzzing algorithm for input testing (zeller2019fuzzing, ; DBLP:conf/ccs/BohmePR16, ) extended for coverage-guided fuzzing with abstract models. The coverage-guided fuzzing algorithm takes the program under test ($S$) and a set of initial test cases ($T_{0}$) as inputs. It maintains the set of test cases to explore ($T$) and the total coverage ($totalCoverage$) of the executed test cases. After each test execution (line 5), the algorithm checks if the execution covers new system behavior (line 6). If the test case covers new behavior, it assigns an energy value to the test case proportional to the new coverage it achieves (line 7) to explore more around the executions that cover more new behavior. The algorithm generates new test cases by mutating such executions, adds them to the set of executions to explore (lines 8-9), and updates the total coverage with the newly covered behavior (line 10). It terminates when a test budget (e.g., specified by a test duration or number of test cases) returns the explored set of test cases together with the test coverage.
@@
Event schedules as test cases. A test case corresponds to a schedule of distributed system events ${e_{0},\ldots,e_{n}}\right>$, where $e_{i}$ is either (i) the delivery of a message or (ii) crashing a process or (iii) restarting it. We collect the set of initial test schedules ($T_{0}$) by randomly scheduling the events in the executions, where we select the next event uniformly at random among the set of all enabled events.
The space of mutations. A difficulty in mutating event schedules is obtaining a feasible schedule after mutating the event order. An offline shuffle of a sequence is likely to produce an infeasible schedule, e.g., the new schedule may not contain all the events that appear in the new execution.
[…]
@@
  • SwapCrashProcesses randomly selects two schedule indices $i,j$ where the recipient of the $\mathtt{buff}_{i}$ is crashed at step $i$ and $\mathtt{buff}_{j}$ is crashed at step $j$, and swaps the positions of the crash events (for schedules with a single crashing process, it changes the process),
  • SwapMaxMessages randomly selects two schedule indices $i,j$ with message delivery events and swaps $q_{i}$ and $q_{j}$.

### v3

@@ Abstract.
We present a coverage-guided testing algorithm for distributed systems implementations. Our main innovation is the use of an abstract formal model of the system that is used to define coverage. Such abstract models are frequently developed in the early phases of protocol design and verification but are infrequently used at testing time. We show that guiding random test generation using model coverage can be effective in covering interesting points in the implementation state space. We have implemented a fuzzer for distributed system implementations and abstract models written in TLA+. Our algorithm achieves better coverage over purely random exploration as well as random exploration guided by different notions of scheduler coverage and mutation. In particular, we show consistently higher coverage on implementations of distributed consensus protocols such as Two-Phase Commit and the Raft implementations in Etcd-raft and RedisRaft and detect bugs faster. Moreover, we discovered 12 previously unknown bugs in their implementations, four of which could only be detected by model-guided fuzzing.
@@
copyright: ccdoi: 10.1145/3763060journalyear: 2025journal: PACMPLjournalvolume: 9journalnumber: OOPSLA2article: 282publicationmonth: 10ccs: Software and its engineering Software testing and debuggingccs: Theory of computation Distributed computing models
@@ 1. Introduction
Coverage-guided fuzzing, which guides test generation toward more coverage, has been effective in exploring diverse executions, mainly in the sequential setting, using structural coverage criteria as a feedback mechanism (Heuse et al., 2022; Zeller et al., 2019). However, adopting coverage-guided fuzzing for testing distributed system implementations is nontrivial since there is no common notion of coverage for distributed system executions. Unfortunately, structural code coverage criteria such as line coverage can ignore the orderings of message interactions in a system, thus missing interesting schedules. On the other hand, more detailed criteria, such as traces of messages, may provide too many coverage goals and thus consider each random trace a new behavior, giving up the advantages of coverage-guided exploration.
@@
In this paper, we propose to use state coverage in an abstract formal model of the system as a coverage criterion and present model-guided fuzzing of distributed systems. Abstract formal models are often developed in the design phase of distributed systems to model and formally analyze the underlying protocols (Lamport, 2002; Desai et al., 2013; Deligiannis et al., 2015, 2021; Bornholt et al., 2021; Newcombe et al., 2015). We show that these artifacts are also beneficial in the continuous testing infrastructure of the implementations themselves. Our experiments show that a formal model can serve as a good “guide” for a random testing engine—this is because the formal model often captures the important scenarios of the protocol, and coverage of states in the model correlates well with coverage of interesting behaviors in the implementation.
@@
Our motivation for using abstract model guidance to generate test executions shares common insights with semantic fuzzing (Padhye et al., 2019) and grammar-based fuzzing approaches (Le et al., 2019; Gopinath et al., 2020) used for testing sequential programs. Semantic fuzzing aims to cover interesting program executions processing program inputs rather than spending exploration budget for exercising uninteresting, syntactic input parsing logic. While a naive fuzzer is likely to generate inputs that cannot pass the input validation and parsing stage, semantic fuzzing generates test inputs that can go deep into the execution. Similarly, a naive event scheduler for distributed systems is likely to produce tests that spend execution budget in exercising uninteresting, network setup stages. For example, it can explore many different orderings of vote messages during the cluster’s leader election phase, barely electing a leader after a prolonged execution. Our approach aims to direct testing toward interesting system behaviors, e.g., processing of user requests once a leader is elected. Similar to grammar-based fuzzing that uses formal specification of the test input to guide test generation, we use an abstract formal model of distributed systems to guide the generation of semantically interesting temporal event schedules.
@@
We have implemented our algorithm for testing distributed systems implementations using TLA+ models (Lamport, 2002) of protocols. In a nutshell, our testing algorithm proceeds as follows. We start by exploring random schedules of messages, but feed the same sequence of messages to the TLA+ model. We use the TLC model checker (Yu et al., 1999) to generate the set of reachable model states corresponding to the explored schedule. We mark a schedule as “interesting” if it covers a new state of the abstract model. We perform, as in coverage-guided fuzzing, a mutation of an interesting schedule by swapping the receipt order of two randomly chosen messages or changing the processes to crash. Applying a mutation to an event schedule gives a new schedule to explore that is similar to the original schedule but likely to exercise new system behavior.
We applied our algorithm to test an implementation of the Two-Phase Commit protocol and two industrial implementations of Raft in Etcd-raft and RedisRaft. Our evaluation shows that model-guided fuzzing leads to higher coverage and can detect bugs faster than pure (unguided) random testing, structural code-coverage guided fuzzing, and trace-based coverage-guided fuzzing. Besides reproducing known bugs, we discovered 12 previously unknown bugs in the implementations of Etcd-raft and RedisRaft. Moreover, four of the new bugs could only be detected by model-guided fuzzing.
@@
    We implement ModelFuzz for testing the two production implementations of the Two-Phase Commit and Raft protocols and evaluate it compared to the existing approaches.
@@
    We discovered 12 unknown bugs in total in the implementations of Etcd-raft and RedisRaft.
@@ 2. Overview
Figure 1 presents an example system implemented in a Coyote-like actor programming framework (Agha, 1985; Deligiannis et al., 2021). The system consists of three processes AppMaster, Worker, and Terminator. Each runs in a separate process and communicates with each other by exchanging asynchronous messages. AppMaster receives client requests, coordinates their execution by a Worker process, and manages the termination of the worker using a Terminator process. It accepts the registration messages from Worker and Terminator and registers them (lines 5-11). Upon receipt of a client request Request(r), it checks whether the cluster is ready by checking the registrations of the worker and terminator. If it is ready, it sends Execute(r) to Worker to handle the request and sends Terminate(w) to Terminator (lines 13-17). The Worker and Terminator register to the AppMaster upon their initialization (lines 26, 39). The Worker handles Execute(r) by processing it (line 30) and Flush by cleaning up its buffers (line 34). The Terminator handles Terminate by sending Flush to the worker (line 42).
@@ 3 StateMachine terminator;
16 terminator.send(Terminate(worker));
@@ 24 Buffer buffer;
The above implementation has a message race bug (Leesatapornwongsa et al., 2016), which occurs in a particular delivery ordering of the asynchronous messages. The worker code mistakenly omits to check if the buffer is null before processing a task (line 31). This causes the worker to access a null pointer while processing the client request if Flush is delivered to it before Execute(r). Figures 1 and 1 illustrate a correct and an incorrect execution of the implementation.
@@
Coverage-guided testing
Coverage-guided techniques guide the generation of test executions toward unexplored system behaviors to increase test coverage and search for bugs more efficiently. These methods track the coverage of each test execution and use this information while generating new test cases. However, structural test coverage metrics, such as line or branch coverage, are ineffective for assessing the coverage of distributed system executions since they do not capture the concurrency behavior of the distributed systems. For example, both correct and buggy executions in Figures 1 and 1 hit exactly the same lines and branches of the program in Figure 1 but with different processing orders of the messages.
@@
Instead, we propose to use an abstract, formal, model of the system (e.g., its TLA+ model (Lamport, 2002)) to assess test coverage and guide the test generation.
@@
We could have alternatively considered trace-based coverage based on the Mazurkiewicz traces arising out of an execution. Mazurkiewicz traces associate two sequences of messages with the same Mazurkiewicz trace if they only reorder independent messages, whose reordering does not affect the system’s behavior. Our state-based notion is coarser than traces while keeping the essential information about the system’s behavior. For example, although the messages Register(w) and Register(t) are dependent, their relative ordering does not affect the system state. On the other hand, the ordering of Request(r) affects the reached system state; the system handles the request only if it is delivered after the two registration messages. The given system has 10 possible message orderings with 8 Mazurkiewicz traces (capturing the commutativity of the Execute and Terminate messages). However, the set of all possible system states can be covered by running fewer executions, e.g., only 2 executions for this example.
@@
Consider an example where the following set of executions has been explored:
@@
Coverage-guided testing selects the executions that hit previously unseen coverage classes, and it generates new test executions by mutating them. Trace-based coverage would label all these executions as interesting since each belongs to a different coverage class (i.e., they deliver Request(r), Register(w), and Register(t) to the same process in a different order). Generating new tests around all of these executions is likely to result in many redundant runs since many of them (E1-E4) already demonstrate the same system behavior.
@@ 3. Coverage-Guided Fuzzing with Abstract Models
We use coverage-guided fuzzing for exploring distributed system executions and guiding the test generation using abstract model coverage. Our method is complementary to the traditional fuzzing methods (e.g., AFL (Zalewski, [n. d.])), which explore the space of program inputs, as ModelFuzz explores the space of event schedules in distributed systems. In this section, we define executions and explain how we adopt the coverage-guided fuzzing approach for the model-guided exploration of distributed system executions.
@@ 3.1. Executions of Distributed Systems
A distributed system S{S} consists of a set of processes that concurrently operate on their local states and communicate by exchanging asynchronous messages. The processes are equipped with message buffers that keep the messages sent to that process, and they process the messages in their buffers serially. Upon handling a message, a process can update its local state and/or send new messages to the processes.  [^2] We consider FIFO message queues that preserve the order of messages between the same sender-receiver pairs, as in distributed system frameworks such as P (Desai et al., 2013), Coyote (Deligiannis et al., 2021), or Akka (Lightbend, Inc, 2023).
Let $\mathtt{Procs}$ be the set of processes and $\mathtt{Msgs}$ be the set of all messages exchanged in the system. We represent the delivery of a message $\mathtt{msg}\in\mathtt{Msgs}$ to its receiver process by an event $e=\left<{\mathtt{recv},\mathtt{send},\mathtt{msg}}\right$> for $\mathtt{recv},\mathtt{send}\in\mathtt{Procs}$, where $\mathtt{recv}(e)$ is the receiver of the message, $\mathtt{send}(e)$ the sender, and $\mathtt{msg}(e)$ is the message. Let $\Sigma$ be the set of all message delivery events. A state $s$ of the system is $s:\langle\mathtt{Buffer}:\mathtt{Procs}\times\mathtt{Procs}\mapsto[\mathtt{Msgs}]\rangle$ where $\mathtt{Buffer}$ is a map from a sender-receiver pair of processes to the list of messages in their message buffers. A transition in the system picks a buffer $\mathtt{buff}$ and executes the first message $\mathtt{msg}$ in it, i.e., running an event $e=\left<{\mathtt{recv},\mathtt{send},\mathtt{msg}}\right$>. Executing the message can lead to the creation of messages $m_{i}$ sent to $\mathtt{recv}$, i.e., $\mathtt{send}$ may send new messages to some processes upon processing $m$. The new state $s^{\prime}$ is obtained by removing $m$ from $s(\mathtt{buff})$ and appending $m_{i}$ to $s(\mathtt{Buffer}(\mathtt{send},\mathtt{recv}))$ for each $i$, and we write $s\xrightarrow{e}s^{\prime}$. An execution of the distributed system is a sequence $s_{0}\xrightarrow{e_{0}}s_{1}\xrightarrow{e_{1}}\ldots\xrightarrow{e_{n}}s_{n+1}$ of states $s_{i}$ and events $e_{i}$. In addition to delivering messages, we introduce two other types of events $\left<{\mathtt{proc},\mathtt{crash}}\right$>, and $\left<{\mathtt{proc},\mathtt{restart}}\right$>, which respectively correspond to crash or restart events of a process $\mathtt{proc}$. We call the sequence $\left<{e_{0},\ldots,e_{n}}\right$> an event schedule.
@@ 3.2. Coverage-guided Fuzzing of Distributed Systems
Output: The set of covered abstract states totalCoveragetotalCoverage
@@ 10          $p\leftarrow\mathtt{assignEnergy}(coverage)$
12             $^{\prime}=\mathtt{mutate}(t) ; TT$$\leftarrow T\cup T^{\prime}$
@@ 21 $states\leftarrow\mathtt{executeAbstractModel}(M,actions)$
Algorithm 1 Coverage-guided fuzzing using abstract models. The statements that differ from traditional fuzzers are highlighted in blue.
Algorithm 1 shows the coverage-guided fuzzing algorithm for generating test inputs (Zeller et al., 2019; Böhme et al., 2016), which we extended for generating test schedules using coverage-guided fuzzing with abstract models.
The coverage-guided fuzzing algorithm takes the program under test ($S$) and a set of initial test cases ($T_{0}$) as inputs. It maintains the set of test cases to explore ($T$) and the total coverage ($totalCoverage$) of the executed test cases. After each test execution (line 5), the algorithm checks if the execution covers new system behavior (line 6). If the test case covers new behavior, it assigns an energy value to the test case proportional to the new coverage it achieves (line 7) to explore more around the executions that cover more new behavior. The algorithm generates new test cases by mutating such executions, adds them to the set of executions to explore (lines 8-9), and updates the total coverage with the newly covered behavior (line 10). It terminates when a test budget (e.g., specified by a test duration or number of test cases) returns the explored set of test cases together with the test coverage.
@@
Event schedules as test cases.
A test case corresponds to a schedule of distributed system events $\left<{e_{0},\ldots,e_{n}}\right$>, where $e_{i}$ is either (i) the delivery of a message or (ii) crashing a process or (iii) restarting it. We collect the set of initial test schedules ($T_{0}$) by randomly scheduling the events in the executions, where we select the next event uniformly at random among the set of all enabled events.
The space of mutations.
Mutations are used to extend the search around interesting schedules that cover new coverage goals. They introduce small changes to an event schedule to produce new executions that are similar to the original schedule but likely to exercise new behavior. A difficulty in mutating event schedules is obtaining a feasible schedule after mutating the event order. An offline shuffle of a sequence is likely to produce an infeasible schedule, e.g., the new schedule may not contain all the events that appear in the new execution.
We represent event schedules by referring to the events using the processes that run them rather than directly referring to the delivered messages. We represent a test case as a sequence of abstract events $s=\left<{\mathtt{buff}_{0}:a_{0}}\right>\ldots\left<{\mathtt{buff}_{n}:a_{n}}\right$> with $a_{0},\ldots,a_{n}\in\{\mathtt{deliver},\mathtt{crash},$ $\mathtt{restart}\}$. Event $\langle\mathtt{buff}_{i}:\mathtt{deliver}\rangle$ delivers the first message in $\mathtt{buff}_{i}$ to its recipient process, $\left<{{\mathtt{buff}_{i}:\mathtt{crash}}}\right$> crashes the recipient process of $\mathtt{buff}_{i}$, and $\left<{{\mathtt{buff}_{i}:\mathtt{restart}}}\right$> restarts it. The test case does not explicitly refer to the messages delivered by the $\mathtt{deliver}$ action (i.e., with certain sender and content) but indirectly refers to them using the message buffers whose messages we deliver. Moreover, we parameterize $\mathtt{deliver}$ with the number of messages to deliver, e.g., $\left<{{\mathtt{buff}_{i}:\mathtt{deliver}(n)}}\right$> corresponds to delivering $n$ messages from $\mathtt{buff}_{i}$ to the recipient process. The indirect representation of the messages helps design mutations that result in feasible event schedules while modifying the order of events. We use the following mutations to generate new schedules:
@@
  • SwapCrashProcesses randomly selects two schedule indices $i,j$ where the recipient of the $\mathtt{buff}_{i}$ is crashed at step $i$ and $\mathtt{buff}_{j}$ is crashed at step $j$, and swaps the positions of the crash events (for schedules with a single crashing process, it changes the process to crash),
  • SwapMaxMessages randomly selects two schedule indices $i,j$ with message delivery events and swaps $q_{i}$ and $q_{j}$, i.e., the number of messages to deliver at these positions.